69) Assume that a company provided the following excerpts of information from its flexible budget performance report:

Actual Results Flexible Budget Planning Budget
Flights (q) 52 ? ?
Expenses:
Wages and salaries ($4,000 + $88.00q) $ 8,495 ? $ 8,400

What is the spending variance for wages and salaries expense?

Multiple Choice

  • $176 U

  • $176 F

  • $81 F

  • $81 U

75) Assume the following:

  • The standard labor rate per hour is $17.20.
  • The standard labor-hours allowed per unit of finished goods is 3 hours.
  • The actual quantity of labor hours worked during the period was 44,000 hours.
  • The total actual direct labor cost for the period was $726,000.
  • The company produced 15,000 units of finished goods during the period.

What is the labor rate variance?

Multiple Choice

  • $28,800 U

  • $28,800 F

  • $30,800 U

  • $30,800 F

76) Assume the following:

  • The standard price per pound is $2.00.
  • The standard quantity of pounds allowed per unit of finished goods is 4 pounds.
  • The actual quantity of materials purchased and used in production is 50,000 pounds.
  • The materials price variance is $9,000 U.
  • The company produced 13,000 units of finished goods during the period.

The actual price per pound of direct materials is closest to:

Multiple Choice

  • $1.68.

  • $2.18.

  • $1.78.

  • $2.08.
Answer #1

69)Correct option is "C "- 81 F

Standard cost for actual output= 4000+ (88*52)= 8576

spending variance for wages and salaries expense =Actual cost -Standard cost for actual output

                          = 8495 - 8576

                          = 81 F (ACTUAL COST IS LESS THAN STANDARD COST ,VARIANCE IS FAVORABLE)

75)Correct option is "D" - 30800F

Labor rate variance =Actual labor cost -[Actual hours*Standard rate per hour)

                   = 726000 - [44000*17.20]

                   = 726000 - 756800

                  = 30800 F

76)COrrect option is "B"-2.18

Material price variance =Actual quantity [Actual rate per pound -standard rate per pound ]

9000 = 50000[Actual rate per pound - 2]

9000/50000= Actual rate per pound-2

Actual rate per pound = .18+2

                     = $ 2.18 per pound